OpenStudy (anonymous):

Solve for x: 2 over 5x + 2 = 10 x = 20 x = 24 x = 4 x = 8

OpenStudy (anonymous):

OHH it's (2/5x) + 2 = 10

OpenStudy (anonymous):

2/5 x = 8

OpenStudy (anonymous):

8 * 5/2 = 20

OpenStudy (anonymous):

x = 20

OpenStudy (anonymous):

20 or 8?

OpenStudy (anonymous):

I refer you to my previous statement.

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

better question is now that Freyes have given you the answer.. do you know how to do a similar problem?

OpenStudy (anonymous):

\[\frac25x+2=10\\2x+10=50\\2x=40\\x=20\]